# Spiking Neural Networks for Radio Frequency Interference Detection in Radio Astronomy

### v1

@@ Discussion
This study demonstrates that SNNs can effectively perform time-series segmentation tasks, specifically RFI detection in radio astronomy. Reformulating RFI detection as a time-series segmentation problem, we leveraged the temporal dynamics inherent to spiking neurons to recognise complex patterns. We explored various spike encoding methods, including latency, rate, delta-modulation, delta-exposure and step-forward encodings to convert ‘visibility’ spectrographic data to spikes. We additionally explored optimising network depth and hidden layer widths alongside the spiking neuron and encoding parameters. Our experiments show that simple feed-forward SNNs of first-order leaky-integrate-and-fire neurons perform well on synthetic datasets, achieving near state-of-the-art results and moderately well on real-world data from the LOFAR instrument, despite increased noise and variability, outperforming equivalent ANNs. To our knowledge, this work is the first to train SNNs on a real radio astronomy observation dataset successfully. Our findings suggest latency with divisive normalisation and step-forward (direct exposure) encoding methods performed particularly well. This indicates that SNNs may tolerate several different spike encoding schemes.
@@
A key contribution of this work is introducing a divisive normalisation-inspired pre-processing step. Our results indicate that this pre-processing can significantly improve detection performance, particularly on the synthetic dataset. However, expanding the normalisation kernel size or incorporating information from additional prior time steps may further improve performance on realistic datasets, where background information is more dynamic.
@@
Despite promising initial results, our work has limitations. Training resources on the LOFAR dataset were equivalent to the synthetic HERA dataset; training for extended durations with larger networks may provide additional performance. Directly scaling this work’s methods is yet to be exhausted. Still, the combination of increased computational requirements in simulating large SNNs and the wider configurability of spiking neurons makes exploring this space particularly challenging. Further developments on both the data-preparation and architectural fronts may provide additional performance. For example, capturing the polarity information from the original observational data, developing more sophisticated architectures that consume entire spectrograms in a multi-headed approach or with built-in de-noising capabilities and exploring training encodings are of future interest. Future work could also focus on integrating more sophisticated neuron models, such as sigma-delta or second-order LiF neurons, decoding the generally noisy output spike-trains in a trainable fashion, or casting other spectrographic segmentation tasks as time-series segmentation, such as oceanography or seismic inversion.
@@
Modern radio astronomy is as much a science based on building precision instruments as high-performance computing. Over multi-decade project horizons, integrating cutting-edge computing methods to find performance and efficiency benefits directly translates into greater scientific impact over the life of a telescope. Despite this work’s promising first steps, challenges remain regarding the operational integration of neuromorphic computing into radio astronomy processing. While access to commercial nascent neuromorphic hardware has never been easier, most platforms focus primarily on image-oriented [29, 30] or narrow-width time-series inputs (relative to the needs of radio astronomy) [31]. Platforms like Intel’s Loihi [29, 32] provide greater flexibility, but the practical performance and efficiency of such systems on RFI detection is yet to be known. Moreover, many focus on edge-computing use cases, where small size, weight and power consumption dominate design considerations. In an observatory environment, the opportunity exists to target physically larger, more powerful neuromorphic systems that still provide energy efficiencies relative to contemporary high-performance computing systems. Moreover, a practical claim of RFI detection techniques suitable for real-time processing demands rigorous performance analysis that may impose strict requirements on exposure times, for example, when considering the time-series segmentation method this article presents. Integrating neuromorphic computing into radio astronomy or other data-intensive sciences could be a prime candidate for hardware-software co-design. Finally, there exists the challenge, not unique to potential SNN-based methods, of integrating machine-learned or automated methods, which can be brittle and difficult to adapt quickly, into operational observatories generally.
@@
In conclusion, our investigation continues to pave the way for applying SNNs and neuromorphic computing in radio astronomy, offering unique advantages over contemporary methods. We contribute to advancing efficient, adaptive data processing techniques essential to the future of radio astronomy while simultaneously finding a new domain to explore the capabilities of neuromorphic computing and SNNs.
@@ Methods
The LOFAR dataset, as fully described by Mesracik et al. [14], consists of 7,500 training spectrograms with corresponding masks produced by AOFlagger [8], along with 109 testing spectrograms that have been expertly hand-labelled. This dataset originates from five observations, totalling 1.7 TB in size. To reduce this volume, the dataset was downsampled by selecting only the first Stokes parameter and randomly sampling 1,500 baselines, resulting in a more manageable size of approximately 10 GB. Combining machine-generated labels and hand-labelled testing examples adds to the complexity and challenge of the dataset in addition to a significantly more difficult noise environment. We trained on 15% of the training samples for our experiments, 1125 spectrograms.
SNN neuron model
This work utilises a first-order leaky integrate and fire neuron model reset with subtraction, found in snnTorch [34]. The linear differential equation modelling the membrane potential $U(t)$ dynamics of a single neuron subject to an input current $I_{in}$ is given as:
@@
All simulations were conducted on the Setonix supercomputer, an HPE Cray EX system located at the Pawsey Supercomputing Centre in Western Australia. Specifically, each trial utilised a single node with a single AMD EPYC 7A53 ‘Trento’ 64-core CPU, eight AMD Instinct MI250X GPUs, and 256 GB of RAM. For SNN network training, we employed the snnTorch framework [34], with distributed training managed by the Lightning library [35] to efficiently leverage the available GPU resources.
@@
We use the tree-structured Parzan estimation algorithm Optuna [36] exposes for multi-variate optimisation. We simultaneously optimised for per-pixel accuracy, Area Under the Receiver Operating Characteristic curve (AUROC), Area Under the Precision-Recall Curve (AUPRC), Accuracy, and F1-Score. AUROC evaluates the ratio of True Positive Rate (TPR) and False Positive Rate (FPR) across all possible classification thresholds. AUPRC gives the ratio of precision and recall across several thresholds, in this case referring to the proportion of correctly classified RFI over all RFI predictions. Recall, in this case, is the TPR. Accuracy is the per-pixel output accuracy. We expect accuracy to be very high for both the synthetic and real radio astronomy datasets as true RFI flags are very sparse; a completely silent network would still be more than 90% accurate. F1-Score is the harmonic mean of precision and recall at a given threshold. The sparsity of RFI in this dataset means we generally expect high AUROC and high accuracy values with lower AUPRC and F1 scores. A good-performing model must be silent most of the time while providing confident RFI detections, hence our inclusion of both raw accuracy and F1-scores. Results for each trial were uploaded to a persistent database from which Optuna read, allowing each trial to execute as an individual job. Each hyper-parameter trial ran for 50 epochs with an initial learning rate of $31e-$ 3. However, utilising Lightning’s plateau scheduler reduced the learning rate by half for every ten epochs with no improvement in validation loss. Batch size for all trials was fixed at $36$. Table 5 contains the ranges and descriptions of each hyper-parameter tuned for each encoding. In Table 5, ‘Num hidden’ refers to the number of neurons in each hidden layer. ‘Num layers’ refers to the number of hidden layers in the network. Beta defines the excitability of LiF neurons within the network, and exposure is a parameter for the encoding method used in all but the delta-modulation method. Beta and exposure hold no meaning for ANN networks and are omitted for these trials. For all encoding methods, we took the trial with the highest-scoring results in the most categories into repeated trials, tie-breaking by F1-Score performance.
Table 5: Parameter ranges for attributes included in the final hyper-parameter searches.
Attribute Parameter Range
Num Hidden 128, 256, 512
Num Layers 2 - 6
Beta 0.5 - 0.99
Exposure 1 - 64
@@
Repeat trial simulations were run for 100 epochs, each with the same learning rate scheduling and batch size as the hyper-parameter trials.
@@
Figure 5(b) contains an example input raster plot.
@@
Figure 5(c) contains an example input raster plot.
@@
Figure 5(d) contains an example input raster plot.
@@
Delta exposure is a hybrid of the latency and delta-modulation methods. Input encoding is similar to delta modulation. However, we include an exposure time to give the network more time to react to the input stimulus. Output decoding is also simplified to the scheme described in equation 7. Figure 5(e) contains an example raster for this scheme.
@@
This exposure mode presents spikes to the network only on the first exposure step for each time step, with silence following. Figure 5(f) contains an example raster plot for this exposure method.
@@
This exposure mode presents spikes to the network at every exposure step for each time step. Figure 5(g) contains an example raster plot for this exposure method.
@@
This exposure mode presents spikes to the network latency encoded at exposure time 00 where a spike is present and exposure time $E$ otherwise. Figure 5(h) contains an example raster plot for this exposure method.
(a) The original input spectrogram and associated RFI mask.
(b) Latency encoding reveals clear banding and blip features.
(c) Rate encoding reveals most features clearly but struggles with isolated blips.
(d) Delta encoding condenses frequency-isolated RFI into a few spikes, making thresholding during encoding critical.
(e) Delta exposure encoding improves detail but still struggles with frequency-isolated RFI.
(f) Step-Forward encoding with ‘first’ exposure mode clearly shows banding but is hard to interpret.
(g) Step-Forward encoding with ‘direct’ exposure is active but shows the difficulty in identifying bands.
(h) Step-Forward encoding example with ‘latency’ exposure effectively identifies bands and blips.
Figure 5: Spike Raster Plots for Various Encodings Applied to Synthetic HERA Data. Encodings are exposed for four time steps where applicable, showcasing how different methods capture distinct RFI features.
@@
This study’s datasets are openly available online [37].
@@
Source code is openly available [38].

### v2

@@ Discussion
This study demonstrates that SNNs can effectively perform time-series segmentation tasks, specifically RFI detection in radio astronomy. Reformulating RFI detection as a time-series segmentation problem, we leveraged the temporal dynamics inherent to spiking neurons to recognise complex patterns. We explored various spike encoding methods, including latency, rate, delta-modulation, delta-exposure and step-forward encodings to convert ‘visibility’ spectrographic data to spikes. We additionally explored optimising network depth and hidden layer widths alongside the spiking neuron and encoding parameters. Our experiments show that simple feed-forward SNNs of first-order leaky integrate and fire (LiF) neurons perform well on synthetic datasets, achieving near state-of-the-art results and SNNs of second-order LiF neurons perform moderately well on real-world data from the LOFAR instrument, despite increased noise and variability, outperforming equivalent ANNs and previous ANN-to-SNN based work. To our knowledge, this work is the first to train SNNs on a real radio astronomy observation dataset successfully. The performance discrepancy between the two neuron types is significant, with the simpler neuron performing well on the simpler dataset, and vice-versa for the real dataset. There is a curious overlap between the more biologically accurate neuron model performing uniquely better on the real dataset, handling noise more adeptly, which the simpler model might not capture as easily. While these results provide an initial baseline for future work to compare against, the significant gap between synthetic and real dataset performance highlights the limits of a directly supervised approach to flagging with SNNs. Our findings suggest latency with divisive normalisation and step-forward (direct exposure) encoding methods performed particularly well. This indicates that SNNs may tolerate several different spike encoding schemes.
@@
A key contribution of this work is introducing a divisive normalisation-inspired pre-processing step. Our results indicate that this pre-processing can significantly improve detection performance, particularly on the synthetic dataset. However, expanding the normalisation kernel size or incorporating information from additional prior time steps may further improve performance on realistic datasets, where background information is more dynamic. With additional investigation into the effects of this pre-processing method, we uncover that this divisive normalisation step has a different impact on the simulated data to the real dataset, where supervision masks are less closely aligned to the underlying RFI. This is turn suggests adapting the lessons learned in training SNNs on realistic radio astronomy data into other formulations of the RFI detection task, such as an anomaly detection task.
@@
Despite promising initial results, our work has limitations. The training procedure on the LOFAR dataset was equivalent to the synthetic HERA dataset, whereas training with different schedules, for extended time, with larger or more complex networks, or even a simulation-to-real approach, may yield additional performance. Directly scaling this work’s methods is yet to be exhausted. Still, the combination of increased computational requirements in simulating large SNNs and the wider configurability of spiking neurons makes exploring this space particularly challenging. Further developments on both the data-preparation and architectural fronts may provide additional performance. For example, capturing the polarity information from the original observational data, developing more sophisticated architectures that consume entire spectrograms in a multi-headed approach, incorporating built-in de-noising capabilities, convolutional feature extraction, spiking LSTM networks and exploring training encodings are all of future interest. For example, SNNs have shown promise in audio classification tasks for instance [27] where the delays between neurons play a significant role in incorporating long-range temporal information. Moreover, to address the difficulty in finding appropriate supervised training data in radio astronomy, finding self-supervised methods [15] and adapting those for direct SNN training would potentially be very powerful. Future work could also focus on integrating more sophisticated neuron models, such as sigma-delta or second-order LiF neurons, decoding the generally noisy output spike-trains in a trainable fashion, or casting other spectrographic segmentation tasks as time-series segmentation, such as oceanography or seismic inversion.
@@
Modern radio astronomy is as much a science based on building precision instruments as high-performance computing. Over multi-decade project horizons, integrating cutting-edge computing methods to find performance and efficiency benefits directly translates into greater scientific impact over the life of a telescope. Despite this work’s promising first steps, significant challenges remain regarding the operational integration of neuromorphic computing into radio astronomy processing. While access to commercial nascent neuromorphic hardware has never been easier, most platforms focus primarily on image-oriented [34, 35] or narrow-width time-series inputs (relative to the needs of radio astronomy) [36]. Platforms like Intel’s Loihi [34, 37] provide greater flexibility, but the practical performance and efficiency of such systems on RFI detection is yet to be known. Moreover, many focus on edge-computing use cases, where small size, weight, and power consumption dominate design considerations. Nevertheless, we discuss preliminary feasibility and efficiency estimates in the methods section, finding RFI detection as time-series segmentation feasible on the order of <50 mW per baseline with contemporary hardware. In an observatory environment, the opportunity exists to target physically larger, more powerful neuromorphic systems that still provide energy efficiencies relative to contemporary high-performance computing systems. Moreover, a practical claim of RFI detection techniques suitable for real-time processing demands rigorous performance analysis that may impose strict requirements on exposure times, for example, when considering the time-series segmentation method this article presents. Integrating neuromorphic computing into radio astronomy or other data-intensive sciences could be a prime candidate for hardware-software co-design. Finally, there exists the challenge, not unique to potential SNN-based methods, of integrating machine-learned or automated methods, which can be brittle and difficult to adapt quickly, into operational observatories generally.
@@
In conclusion, our investigation continues to pave the way for applying SNNs and neuromorphic computing in radio astronomy, potentially offering unique advantages over contemporary methods. We contribute to advancing efficient, adaptive data processing techniques essential to the future of radio astronomy while simultaneously finding a new domain to explore the capabilities of neuromorphic computing and SNNs.
@@ Methods
The LOFAR dataset, as fully described by Mesracik et al. [15], consists of 7,500 training spectrograms with corresponding masks produced by AOFlagger [9], along with 109 testing spectrograms that have been expertly hand-labelled. This dataset originates from five observations, totalling 1.7 TB in size. To reduce this volume, the dataset was downsampled by selecting only the first Stokes parameter and randomly sampling 1,500 baselines, resulting in a more manageable size of approximately 10 GB. Combining machine-generated labels and hand-labelled testing examples adds to the complexity and challenge of the dataset in addition to a significantly more difficult noise environment. We trained on 15% of the training samples for our experiments, 1125 spectrograms. The motivation for selecting both this dataset and the HERA simulation dataset is that other methods have been benchmarked on these datasets, which provides a convenient way to compare the performance of our methods against traditional algorithmic approaches and other machine learning based RFI detection schemes.
SNN Neuron Models
This work utilises two neuron types, a first-order leaky integrate and fire neuron model reset with subtraction, and a second-order leaky integrate and fire model, both found in snnTorch [39]. For the first-order model The linear differential equation modelling the membrane potential $U(t)$ dynamics of a single neuron subject to an input current $I_{in}$ is given as:
@@
A second-order LiF model adds an additional decay term, modelling both a membrane and synaptic state, replacing $I_{in}$ with $I_{syn}$ modelled as:
$I_{syn}[t+1]=\alpha I_{syn}[t]+I_{in}[t+1]$ (6)
which adds an additional parameter $\alpha$ which we initialize to the same values as $\beta$, since both are adjusted during training. The second stateful element allows in principle for learning more nuanced temporal relationships, but is also more difficult to train. We provide results for both datasets using both neuron models, reporting on the best results in the main text and adding the remaining results in Supplementary Table 1 for the HERA dataset and Supplementary Table 2 for the LOFAR dataset.
@@
All simulations were conducted on the Setonix supercomputer, an HPE Cray EX system located at the Pawsey Supercomputing Centre in Western Australia. Specifically, each trial utilised a single node with a single AMD EPYC 7A53 ‘Trento’ 64-core CPU, eight AMD Instinct MI250X GPUs, and 256 GB of RAM. For SNN network training, we employed the snnTorch framework [39] (v0.9.0), with distributed training managed by the Lightning library [40] (v2.2.0) to efficiently leverage the available GPU resources.
@@
We use the tree-structured Parzan estimation algorithm Optuna [41] exposes for multi-variate optimisation. We simultaneously optimised for per-pixel accuracy, Area Under the Receiver Operating Characteristic curve (AUROC), Area Under the Precision-Recall Curve (AUPRC), Accuracy, and F1-Score. AUROC evaluates the ratio of True Positive Rate (TPR) and False Positive Rate (FPR) across all possible classification thresholds. AUPRC gives the ratio of precision and recall across several thresholds, in this case referring to the proportion of correctly classified RFI over all RFI predictions. Recall, in this case, is the TPR. Accuracy is the per-pixel output accuracy. We expect accuracy to be very high for both the synthetic and real radio astronomy datasets as true RFI flags are very sparse; a completely silent network would still be more than 90% accurate. F1-Score is the harmonic mean of precision and recall at a given threshold. The sparsity of RFI in this dataset means we generally expect high AUROC and high accuracy values with lower AUPRC and F1 scores. A good-performing model must be silent most of the time while providing confident RFI detections, hence our inclusion of both raw accuracy and F1-scores. Results for each trial were uploaded to a persistent database from which Optuna read, allowing each trial to execute as an individual job. Each hyper-parameter trial ran for 50 epochs with an initial learning rate of 1e31e-3. However, utilising Lightning’s plateau scheduler reduced the learning rate by half for every ten epochs with no improvement in validation loss. Batch size for all trials was fixed at $36$. Supplementary Table 3 contains the ranges and descriptions of each hyper-parameter tuned for each encoding. For all encoding methods, we took the trial with the highest-scoring results in the most categories into repeated trials, tie-breaking by F1-Score performance.
@@
Repeat trial simulations were run for 100 epochs, each with the same learning rate scheduling and batch size as the hyper-parameter trials. Since the HERA and LOFAR datasets are of different sizes, however, despite the same number of epochs, the LOFAR trials are subject to more gradient updates per epoch, and due to the longer exposure times, the SNNs are evolved for a longer time proportional to this exposure parameter.
@@
Figure 5.b contains an example input raster plot.
@@
Figure 5.c contains an example input raster plot.
@@
Figure 5.d contains an example input raster plot.
@@
Delta exposure is a hybrid of the latency and delta-modulation methods. Input encoding is similar to delta modulation. However, we include an exposure time to give the network more time to react to the input stimulus. Output decoding is also simplified to the scheme described in equation 8. Figure 5.e contains an example raster for this scheme.
@@
This exposure mode presents spikes to the network only on the first exposure step for each time step, with silence following. Figure 5.f contains an example raster plot for this exposure method.
@@
This exposure mode presents spikes to the network at every exposure step for each time step. Figure 5.g contains an example raster plot for this exposure method.
@@
This exposure mode presents spikes to the network latency encoded at exposure time 0 where a spike is present and exposure time $E$ otherwise. Figure 5.h contains an example raster plot for this exposure method.
Figure 5: Spike Raster Plots for Various Encodings Applied to Synthetic HERA Data. Encodings are exposed for four time steps where applicable, showcasing how different methods capture distinct Radio Frequency Interference (RFI) features. Sub-panels depict: (a) original input spectrogram and associated RFI mask; (b) latency encoding reveals clear banding and blip features; (c) rate encoding reveals most features clearly but struggles with isolated blips; (d) delta encoding condenses frequency-isolated RFI into a few spikes, making thresholding during encoding critical; (e) delta exposure encoding improves detail but still struggles with frequency-isolated RFI; (f) step-forward encoding with ‘first’ exposure mode clearly shows banding but is hard to interpret; (g) step-forward encoding with ‘direct’ exposure is active but shows the difficulty in identifying bands, and (h) step-forward encoding example with ‘latency’ exposure effectively identifies bands and blips.
Dataset Preprocessing
Here we discuss our preprocessing methods in more detail and highlight the differences between the simulated HERA and observed LOFAR datasets. Specifically, by highlighting these differences, we aim to at least partially explain the discrepancy in task performance of SNNs trained to perform RFI detection in a supervised manner. Aside from divisive normalisation, we preprocess the HERA dataset by performing min-max scaling between values one standard deviation below the mean and four standard deviations above the mean. We preprocess the LOFAR dataset by performing min-max scaling between values three standard deviations below the mean and 95 standard deviations above the mean. Both datasets are scaled in this manner before applying divisive normalisation. We chose these methods and values based on prior literature [15].
Table 7 contains the maximum, mean, and median values of the pixels classified as noisy and noiseless in each dataset, both with and without divisive normalisation and under an older and revised pre-processing pipeline for the LOFAR dataset.
Table 7: Summary statistics between noisy and noiseless pixels in the HERA and LOFAR datasets with and without Divisive Normalisation (DN).
Dataset
Max
Max
Difference
Mean
Mean
Difference
Median
Median
Difference
HERA - Train 1.00 0.8955 0.1045 0.5423 1.00 0.4577 0.5904 0.5904 0.00
HERA - Test 1.00 0.8742 0.1258 0.6616 0.5349 0.1267 0.7221 0.5963 0.1258
HERA - DN - Train 1.00 0.1179 0.8821 0.2670 0.00 0.2670 0.2370 0.00 0.2370
HERA - DN - Test 1.00 0.0996 0.9004 0.2579 0.00 0.2579 0.2207 0.00 0.2207
LOFAR - Train 1.00 0.7612 0.2388 0.2226 0.0434 0.1792 0.1424 0.00 0.1424
LOFAR - Test 1.00 1.00 0.00 0.3926 0.0087 0.3839 0.3039 0.00 0.3039
LOFAR - DN - Train 0.9865 0.9865 0.00 0.0623 0.0105 0.0518 0.00 0.00 0.00
LOFAR - DN - Test 0.9921 0.9921 0.00 0.1187 0.0068 0.1119 0.00 0.00 0.00
We see that in the HERA dataset there are significant differences in the maximum values in the noiseless and noisy supervision masks which are extended by the divisive normalisation technique. Whereas in the LOFAR dataset, this is not the case, in fact, there is no difference at all. In the LOFAR dataset, supervision masks are provided as bounding areas which are not as tightly bound to the RFI as in the simulated, and therefore perfect, supervision masks. Moreover, the differences between the train and test sets for the HERA dataset are generally closer than those found in the LOFAR dataset which suggests that the behaviours learned in the training set may apply to the test set more readily than those discovered on the LOFAR training set.
Patching and Banding
Patching, where we split each spectrogram into $32\times 32$ sized sections for individual inference is necessary to make training practical under backpropagation through time. This approach is a well known approach to dealing with radio astronomy spectrographic data [15] and while convenient, this does mean the same SNN model is exposed to all frequency channels during training. In some trials the model learns a bias to flag particular (relative) frequency inputs or time-steps consistently, leading to visible banding in the output. Once reconstructed, this appears as regular banding across an entire spectrogram. Another cause of a similar error is a quirk of the latency encoding and decoding method. When decoding specifically, any output (for a single spectrogram time-step) over an exposure time occurring before the final time-step is considered RFI, and the model is trained to output spikes at the last time-step to indicate no RFI (having spikes on all outputs is necessary to allow backpropagation to be effective, otherwise the neurons ‘die’ and cease responding to any inputs). Where the network lacks confidence, it can output spikes too early which contributes to the consistent lines across patch boundaries. Supplementary Figure 1 contains example inference of this phenomena. To encourage the SNNs to avoid learning aggressive biases, hyper-parameters such as neuron thresholds and decay rates, in addition to learning rates helps ameliorate this issue. Ideally, a separate model would be trained on a consistent section of the frequency spectrum.
Computational and Energy Efficiency
While not the main focus of this work, we remark briefly on the potential computational and energy efficiency of our proposed method to provide intuitive operational motivation behind exploring SNNs for RFI detection. Specifically, based on available spec-sheet data, we provide a lower and upper estimate on power consumption per baseline for SynSense Xylo hardware [36]. We additionally reason about the clock speeds required for real-time processing of RFI under our time-series segmentation approach, which has very different ergonomics to image-based processing solutions. Finally, we compare the number of parameters in our models to those of other state-of-the-art solutions, to provide intuition about what kind of practical difference a move to neuromorphic computing could make.
[…]
Regarding real-time suitability, RFI detection would need to occur within the integration time of the instrument itself; processing a spectrographic time-step in $\approx$ 3.52 seconds (in the case of the HERA simulation) [43]. Even with 64 exposure time-steps, the maximum we consider in this work, and a very conservative assumption of one million operations per exposure, this places a minimum clock speed of 32MHz, which is significantly lower than the listed clock rate of 50MHz [42]. Again, the point here is not to make a concrete claim around actual real-time performance, but to indicate that a follow-up study with real neuromorphic hardware is deserved and likely to yield positive results.
Finally, to remark on model size, our HERA models contain 73.7k and 203k parameters for the original and normalised datasets, respectively, while our LOFAR models contain 32.8k and 16.4k parameters, respectively. The counter-intuitive observation that the models for the more complex dataset are smaller highlights another challenge in utilising SNNs; supervised training requires error to propagate through time. Where the supervising signal is sparse, as in the LOFAR dataset, it makes learning difficult; however, each parameter in a smaller network will receive relatively more supervision signal in each pass. Even our largest models, compared to recent studies into the computational requirements of deep learning methods in RFI detection [18], where UNet-style models use up to 800k parameters, are significantly smaller. This difference highlights the key driving benefit of moving to a time-series approach to RFI detection with SNNs; the input data rate is constrained to the number of instrument channels, and the time axis is pushed into the network itself.
While further work is required to properly characterise the computational and energy efficiency of a neuromorphic approach to RFI detection, this work provides a precedent for further investigation. It suggests that such a method could provide significant operational benefits.
@@
This study’s original datasets are openly available online [43] under a CC BY 4.0 license.
@@
Source code is openly available [44] under an MIT license.